# AnoFPDM: Anomaly Detection with Forward Process of Diffusion Models for Brain MRI

### v3

@@ Abstract
Weakly-supervised diffusion models in anomaly segmentation, which leverage image-level labels and bypass the need for pixel-level labels during training, have shown superior performance over unsupervised methods, offering a cost-effective alternative. Traditional methods that rely on iterative image reconstruction are not fully weakly-supervised due to their dependence on costly pixel-level labels for hyperparameter tuning in inference. To address this issue, we introduce Anomaly Detection with Forward Process of Diffusion Models (AnoFPDM), a fully weakly-supervised framework that operates without image reconstruction and eliminates the need for pixel-level labels in hyperparameter tuning. By using the unguided forward process as a reference, AnoFPDM dynamically selects hyperparameters such as noise scale and segmentation threshold for each input. We improve anomaly segmentation by aggregating anomaly maps from each step of the guided forward process, which strengthens the signal of anomalous regions in the aggregated anomaly map. Our framework demonstrates competitive performance on the BraTS21 and ATLAS v2.0 datasets, showcasing its effectiveness compared to state-of-the-art methods. Code is available at ***.
@@ 1 Introduction
Figure 1: The diagram of the dynamical selection and SAMs aggregation. Given the selected guidance strength $w$, the noise scale $t_{e}$ and threshold $Q^{*}$ are dynamically selected for each input $\boldsymbol{x}_{0}$ according to the divergence $M_{t_{e}}$ between HFP and UFP calculated in each forward step. The SAMs are aggregated to enhance the signal strength of anomalous regions.
@@ 3.1 Diffusion Models
The variation DDIM [19] is non-Markovian, incorporating the input $\boldsymbol{x}_{0}$ into the forward process. The sampling process of DDIM is factorized as $q(\boldsymbol{x}_{1:T}|\boldsymbol{x}_{0})=q(\boldsymbol{x}_{T}|\boldsymbol{x}_{0})\prod_{t=2}^{T}q(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t},\boldsymbol{x}_{0})$. The forward transition from $\boldsymbol{x}_{t-1}$ to $\boldsymbol{x}_{t}$ can be obtained by Bay’s rule with the closed form $q(\boldsymbol{x}_{t+1}|\boldsymbol{x}_{t},\boldsymbol{x}_{0})=\mathcal{N}\left(\sqrt{\bar{\alpha}_{t+1}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t+1}-\sigma_{t}^{2}}\frac{\boldsymbol{x}_{t}-\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}}{\sqrt{1-\bar{\alpha}_{t}}},\sigma_{t}^{2}\boldsymbol{I}\right)$ for $t\geq 2$, where $\sigma_{t}$ is set to 0 to enable the deterministic perturbation. Then, we can perturb inputs iteratively by
@@ 4.2 Fixed guidance strength selection
We determine the optimal guidance strength $w^{*}$ through grid search. On the validation set, we compute the cosine similarity for each sample for each candidate guidance strength $w$ to establish thresholds $Cos_{w}$ that yields best classification accuracy. On the testing set without image-level labels, we use the threshold $Cos_{w^{*}}$ corresponding to the selected guidance strength $w^{*}$ to classify inputs. Only samples classified as unhealthy undergo further segmentation, while those predicted as healthy are assigned all-zero masks for performance evaluation purposes. A higher guidance strength $w$ typically results in better segmentation performance for unhealthy samples but may lead to poorer performance when considering both healthy and unhealthy samples due to misclassification. Therefore, our selection strategy aims to strike a balance by choosing the largest guidance strength that does not significantly reduce accuracy compared to the best accuracy obtained with a smaller candidate. This search typically starts with a small guidance strength, which is incrementally increased to find the optimal value $w^{*}$.
@@ 4.3 Dynamical hyperparameter selection
To segment the anomalous regions from the AAM, we need to find a threshold $Q^{*}$ such that the predicted pixel-level anomalous labels is obtained as $\boldsymbol{H}\geq Q^{*}$. We note that the size of the anomalous regions in the input $\boldsymbol{x}_{0}$ is roughly proportional to the maximum value of $M_{t_{e}}$. Hence, the value $M_{t_{e}}$ serves as a rough indication of the size of the anomalous regions. A smaller quantile of $\boldsymbol{H}$ is selected for a larger anomalous region to include more possible pixels since the signal strength of anomalous regions may not be consistent in $\boldsymbol{H}$, e.g., the signal strength of central anomalous regions is stronger than the edge due to the proposed aggregation. We select the segmentation threshold $Q^{*}$ for each predicted unhealthy input $\boldsymbol{x}_{0}$ by Algorithm 1. The input $M_{max}$, the largest $M_{t_{e}}$ of all validation samples, is obtained from the validation set for scaling and is fixed for the testing set. In our case, we set $a=0.90$, $b=0.98$ for the range of quantile.
@@ 5.3 Main results
For the proposed method, we present results for four setups: (i) AnoFPDM (DDPM) (stochastic encoding): Eq. 2 is used to perturb inputs; (ii) AnoFPDM (DDIM) (deterministic encoding): the inputs are noised by Eq. 4; (iii) AnoFPDM (DDIM $\boldsymbol{S}_{t_{e}}$): similar to AnoFPDM (DDIM), but segmentation is performed using only the SAM at $t_{e}$; (iv) AnoFPDM (DDPM $\boldsymbol{S}_{t_{e}}$): follows the AnoFPDM (DDPM) setup, with segmentation using only the SAM at $t_{e}$. For comparative analysis, we report the results from AnoDDPM [25] with simplex noise and Gaussian noise for fair comparison, DDIM with classifier guidance [24], DDIM with classifier-free guidance [17] and classifier guided conditional diffusion model (CG-CDM) [9]. The first two methods are only trained on healthy samples. The hyperparameters for all comparison methods are optimized using a grid search on 1,000 mixed samples with pixel-level labels for both setups. Notably, our methods do not require pixel-level labels for tuning. The postprocessing for segmentation is detailed in Appendix E.
Our method, especially AnoFPDM (DDIM) with deterministic encoding, demonstrates superior performance in terms of DICE and IoU scores, particularly in scenarios only involving unhealthy samples. AnoDDPM with simplex noise achieves AUPRC scores comparable to ours. However, its segmentation performance in unhealthy settings is hindered by a fixed threshold. Across both datasets, our approach consistently balances performance between mixed and unhealthy setups, notably excelling in the ATLAS v2.0 dataset where other comparison methods struggle to detect anomalous regions. We discuss further variations in performance across both datasets in Appendix F. The qualitative results, displayed in Fig. 4, underscore our method’s ability to enhance the signal strength of anomalous regions, which is confirmed by higher AUPRC scores. Notably, the signal strength at the edges of anomalous regions is weaker compared to their central parts. This effect becomes more pronounced in larger anomalous regions. More qualitative results are exhibited in Appendix D.
@@ 6 Conclusion
We propose a novel weakly-supervised method, AnoFPDM, for anomaly detection. This method dynamically selects the noise scale and threshold for each input while maintaining a fixed guidance strength selection. In quantitative evaluations, our method surpasses previous approaches and demonstrates enhanced signal strength of anomalous regions in qualitative assessments. Unlike traditional methods, our selection process does not involve pixel-level labels, making it more practical for real-world applications.

### v4

@@ Abstract
Weakly-supervised diffusion models in anomaly segmentation, which leverage image-level labels and bypass the need for pixel-level labels during training, have shown superior performance over unsupervised methods, offering a cost-effective alternative. Traditional methods that rely on iterative image reconstruction are not fully weakly-supervised due to their dependence on costly pixel-level labels for hyperparameter tuning in inference. To address this issue, we introduce Anomaly Detection with Forward Process of Diffusion Models (AnoFPDM), a fully weakly-supervised framework that operates without image reconstruction and eliminates the need for pixel-level labels in hyperparameter tuning. By using the unguided forward process as a reference, AnoFPDM dynamically selects hyperparameters such as noise scale and segmentation threshold for each input. We improve anomaly segmentation by aggregating anomaly maps from each step of the guided forward process, which strengthens the signal of anomalous regions in the aggregated anomaly map. Our framework demonstrates competitive performance on the BraTS21 and ATLAS v2.0 datasets. Code is available at https://github.com/SoloChe/AnoFPDM.
@@ 1 Introduction
Figure 1: The diagram of the proposed method. Given the selected fixed guidance strength $w$, the noise scale $t_{e}$ and threshold $Q^{*}$ are dynamically selected for each input $\boldsymbol{x}_{0}$ according to the divergence $M_{t_{e}}$ between HFP and UFP calculated in each forward step. $\tilde{\boldsymbol{x}}_{0,t}^{h}$ and $\tilde{\boldsymbol{x}}_{0,t}^{\emptyset}$ are HGP and UGP respectively. The SAMs are aggregated to enhance the signal strength of anomalous regions.
@@ 3.1 Diffusion Models
The variation DDIM [20] is non-Markovian, incorporating the input $\boldsymbol{x}_{0}$ into the forward process. The sampling process of DDIM is factorized as $q(\boldsymbol{x}_{1:T}|\boldsymbol{x}_{0})=q(\boldsymbol{x}_{T}|\boldsymbol{x}_{0})\prod_{t=2}^{T}q(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t},\boldsymbol{x}_{0})$. The forward transition from $\boldsymbol{x}_{t-1}$ to $\boldsymbol{x}_{t}$ can be obtained by Bayes’s rule with the closed form $q(\boldsymbol{x}_{t+1}|\boldsymbol{x}_{t},\boldsymbol{x}_{0})=\mathcal{N}\left(\sqrt{\bar{\alpha}_{t+1}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t+1}-\sigma_{t}^{2}}\frac{\boldsymbol{x}_{t}-\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}}{\sqrt{1-\bar{\alpha}_{t}}},\sigma_{t}^{2}\boldsymbol{I}\right)$ for $t\geq 2$, where $\sigma_{t}$ is set to 0 to enable the deterministic perturbation. Then, we can perturb inputs iteratively by
@@ 4.2 Fixed guidance strength selection
We determine the optimal guidance strength $w^{*}$ through grid search. On the validation set, we compute the cosine similarity for each sample for each candidate guidance strength $w$ to establish thresholds $Cos_{w}$ that yields best classification accuracy. On the testing set without image-level labels, we use the threshold $Cos_{w^{*}}$ corresponding to the selected guidance strength $w^{*}$ to classify inputs. Only samples classified as unhealthy undergo further segmentation, while those predicted as healthy are assigned all-zero masks for performance evaluation purposes. A higher guidance strength $w$ typically results in better segmentation performance for unhealthy samples but may lead to poorer performance when considering both healthy and unhealthy samples due to misclassification. Therefore, our selection strategy aims to strike a balance by choosing the largest guidance strength that does not significantly reduce accuracy compared to the best accuracy obtained with a smaller candidate. This search typically starts with a small guidance strength, which is incrementally increased to find the optimal value $w^{*}$. The selection process is detailed in Algorithm 1, as presented in Appendix C.
@@ 4.3 Dynamical hyperparameter selection
To segment the anomalous regions from the AAM, we need to find a threshold $Q^{*}$ such that the predicted pixel-level anomalous labels is obtained as $\boldsymbol{H}\geq Q^{*}$. We note that the size of the anomalous regions in the input $\boldsymbol{x}_{0}$ is roughly proportional to the maximum value of $M_{t_{e}}$. Hence, the value $M_{t_{e}}$ serves as a rough indication of the size of the anomalous regions. A smaller quantile of $\boldsymbol{H}$ is selected for a larger anomalous region to include more possible pixels since the signal strength of anomalous regions may not be consistent in $\boldsymbol{H}$, e.g., the signal strength of central anomalous regions is stronger than the edge due to the proposed aggregation. We select the segmentation threshold $Q^{*}$ for each predicted unhealthy input $\boldsymbol{x}_{0}$ by Algorithm 1. The input $M_{max}$, the largest $M_{t_{e}}$ of all validation samples, is obtained from the validation set for scaling and is fixed for the testing set. In our case, we set $a=0.90$, $b=0.98$ for the range of quantile. The whole segmentation process is detailed in Algorithm 2, as presented in Appendix C.
@@ 5.3 Main results
For the proposed method, we present results for four setups: (i) AnoFPDM (DDPM) (stochastic encoding): Eq. 2 is used to perturb inputs; (ii) AnoFPDM (DDIM) (deterministic encoding): the inputs are noised by Eq. 4; (iii) AnoFPDM (DDIM $\boldsymbol{S}_{t_{e}}$): similar to AnoFPDM (DDIM), but segmentation is performed using only the SAM at $t_{e}$; (iv) AnoFPDM (DDPM $\boldsymbol{S}_{t_{e}}$): follows the AnoFPDM (DDPM) setup, with segmentation using only the SAM at $t_{e}$. For comparative analysis, we report the results from AnoDDPM [26] with simplex noise and Gaussian noise for fair comparison, DDIM with classifier guidance [25], DDIM with classifier-free guidance [18] and classifier guided conditional diffusion model (CG-CDM) [9]. The first two methods are only trained on healthy samples. The hyperparameters for all comparison methods are optimized using a grid search on 1,000 mixed samples with pixel-level labels for both setups. Notably, our methods do not require pixel-level labels for tuning. The postprocessing for segmentation is detailed in Appendix F.
Our method, especially AnoFPDM (DDIM) with deterministic encoding, demonstrates superior performance in terms of DICE and IoU scores, particularly in scenarios only involving unhealthy samples. AnoDDPM with simplex noise achieves AUPRC scores comparable to ours. However, its segmentation performance in unhealthy settings is hindered by a fixed threshold. Across both datasets, our approach consistently balances performance between mixed and unhealthy setups, notably excelling in the ATLAS v2.0 dataset where other comparison methods struggle to detect anomalous regions. We discuss further variations in performance across both datasets in Appendix G. The qualitative results, displayed in Fig. 4, underscore our method’s ability to enhance the signal strength of anomalous regions, which is confirmed by higher AUPRC scores. Notably, the signal strength at the edges of anomalous regions is weaker compared to their central parts. This effect becomes more pronounced in larger anomalous regions. More qualitative results are exhibited in Appendix E.
@@ 6 Conclusion
We propose a novel weakly-supervised method, AnoFPDM, for anomaly detection. This method dynamically selects the noise scale and threshold for each input while maintaining a fixed guidance strength selection. In quantitative evaluations, our method surpasses previous approaches and demonstrates enhanced signal strength of anomalous regions in qualitative assessments. Unlike traditional methods, our selection process does not involve pixel-level labels, making it more practical for real-world applications. We also discuss the limitations in Appendix G.
Appendices
Appendix A Derivation
We provide the derivation of Eq. (8), (9), (13) and (15) in the main text. For Eq. (8), we have
$\tilde{\boldsymbol{x}}_{0,t}^{h}$ $=\frac{\boldsymbol{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\left[(1+w)\boldsymbol{\epsilon_{\theta}}(\boldsymbol{x}_{t},t,h)-w\boldsymbol{\epsilon_{\theta}}(\boldsymbol{x}_{t},t,\emptyset)\right]}{\sqrt{\bar{\alpha}_{t}}}$
$=\frac{\boldsymbol{x}_{t}}{\sqrt{\bar{\alpha}_{t}}}-A_{t}\left[(1+w)\boldsymbol{\epsilon_{\theta}}(\boldsymbol{x}_{t},t,h)-w\boldsymbol{\epsilon_{\theta}}(\boldsymbol{x}_{t},t,\emptyset)\right]$
$=\frac{\boldsymbol{x}_{t}}{\sqrt{\bar{\alpha}_{t}}}+B_{t}\left[(1+w)\boldsymbol{s_{\theta}}(\boldsymbol{x}_{t},t,h)-w\boldsymbol{s_{\theta}}(\boldsymbol{x}_{t},t,\emptyset)\right]$
$=\frac{\boldsymbol{x}_{t}}{\sqrt{\bar{\alpha}_{t}}}+B_{t}\left[(1+w)\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(\boldsymbol{x}_{t}\mid h\right)-w\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(\boldsymbol{x}_{t}\right)\right]$
$=\frac{\boldsymbol{x}_{t}}{\sqrt{\bar{\alpha}_{t}}}+B_{t}\left[(1+w)\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(h\mid\boldsymbol{x}_{t}\right)+\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(\boldsymbol{x}_{t}\right)\right]$
$=\frac{\boldsymbol{x}_{t}}{\sqrt{\bar{\alpha}_{t}}}+B_{t}\nabla_{\boldsymbol{x}_{t}}\log\left[p_{\theta}(\boldsymbol{x}_{t})p_{\theta}\left(h\mid\boldsymbol{x}_{t}\right)^{1+w}\right]$
$=\frac{\boldsymbol{x}_{t}}{\sqrt{\bar{\alpha}_{t}}}+B_{t}\nabla_{\boldsymbol{x}_{t}}\log\tilde{p}_{\theta}\left(\boldsymbol{x}_{t}\mid h\right)$
where $A_{t}=\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}$ and $B_{t}=\frac{1-\bar{\alpha}_{t}}{\sqrt{\bar{\alpha}_{t}}}$. The score functions are defined as $\boldsymbol{s_{\theta}}(\boldsymbol{x}_{t},t,h)=\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(\boldsymbol{x}_{t}\mid h\right)$ and $\boldsymbol{s_{\theta}}(\boldsymbol{x}_{t},t,\emptyset)=\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(\boldsymbol{x}_{t}\right)$. For Eq. (9), we substitute $\boldsymbol{x}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}$ into Eq. (8).
For Eq. (13) and (15), we have
$M_{t}$ $=\frac{\left\lVert\tilde{\boldsymbol{x}}_{0,t}^{h}-\tilde{\boldsymbol{x}}_{0,t}^{\emptyset}\right\rVert_{2}^{2}}{d}$
$=\frac{A_{t}^{2}(1+w)^{2}}{d}\left\lVert\boldsymbol{\epsilon_{\theta,t}^{h}}-\boldsymbol{\epsilon_{\theta,t}^{\emptyset}}\right\rVert_{2}^{2}$
$=\frac{B_{t}^{2}(1+w)^{2}}{d}\left\lVert\boldsymbol{s^{h}_{\theta,t}}-\boldsymbol{s_{\theta,t}^{\emptyset}}\right\rVert_{2}^{2}$
$=\frac{C_{t}^{2}}{d}\left\lVert\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(\boldsymbol{x}_{t}\mid h\right)-\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(\boldsymbol{x}_{t}\right)\right\rVert_{2}^{2}$
$=\frac{C_{t}^{2}}{d}\left\lVert\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(h\mid\boldsymbol{x}_{t}\right)\right\rVert_{2}^{2}$
where $C_{t}=B_{t}(1+w)$
$\boldsymbol{e}_{t}^{h}$ $=\tilde{\boldsymbol{x}}_{0,t}^{h}-\boldsymbol{x}_{0}$
$=B_{t}\left[(1+w)\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(\boldsymbol{x}_{t}\mid h\right)-w\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(\boldsymbol{x}_{t}\right)-\nabla_{\boldsymbol{x}_{t}}\log q(\boldsymbol{x}_{t})\right]$
$=B_{t}\left((1+w)\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(h\mid\boldsymbol{x}_{t}\right)+\Delta\boldsymbol{s}_{t}\right)$
Appendix B Data Preprocessing and Training Hyperparameters
We standardized preprocessing for both the BraTS21 and ATLAS v2.0 datasets. Each subject was normalized by dividing it by the 99th percentile intensity of foreground voxels, and pixel values were then scaled to the range of $\left[-1,1\right]$. All samples are interpolated to $128\times 128$.
The backbone U-net is adopted from the previous work [18]. Our model is trained on 2 Nvidia A100 GPUs with 80GB memory. The training hyperparameters are summarized in Tab. 2, and we used the same hyperparameters for both dataset.
Diffusion steps 1000
Noise schedule linear
Channels 128
Heads 2
Attention resolution 32,16,8
Channel multiplier 1, 1, 2, 3, 4
Dropout 0.1
EMA rate 0.9999
Optimiser AdamW
Learning rate $1e^{-4}$
$\beta_{1}$, $\beta_{2}$ 0.9, 0.999
Global batch size 64
Null label ratio 0.1
dropout 0.1
Table 2: Training hyperparameters used in our method.
Appendix C Fixed Guidance Selection and Segmentation
We illustrate the fixed guidance selection in Algorithm 2 and outline the complete segmentation process in Algorithm 3.
Input: $n$ sorted candidates $[w_{1},...,w_{n}]$, validation set with image-level labels
for each candidate $w_{i}$:
 calculate cosine similarity for each sample in validation set with Eq. 18
 classify each samples in validation set with cosine similarity threshold $Cos_{w_{i}}$
 get the maximal classification accuracy $Acc_{w_{i}}$ using the optimal threshold $Cos^{*}_{w_{i}}$
end for
$w=\arg\max_{w}Acc_{w_{i}}$
$w^{*}=w_{i}>w$ with $\frac{Acc_{w_{i}}}{Acc_{w}}\approx 0.98$
$Cos_{w^{*}}=Cos^{*}_{w_{i}}$ corresponding threshold
Return $w^{*}$, $Cos_{w^{*}}$
Algorithm 2 Selection of fixed guidance $w^{*}$
Input: fixed guidance $w^{*}$, input $\boldsymbol{x}_{0}$
for each time step $t$
$\tilde{\boldsymbol{x}}_{0,t}^{h}=\frac{\boldsymbol{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\left[(1+w^{*})\boldsymbol{\epsilon_{\theta}}(\boldsymbol{x}_{t},t,h)-w^{*}\boldsymbol{\epsilon_{\theta}}(\boldsymbol{x}_{t},t,\emptyset)\right]}{\sqrt{\bar{\alpha}_{t}}}$ Healthy guided prediction Eq. 7
$\tilde{\boldsymbol{x}}_{0,t}^{\emptyset}=\frac{\boldsymbol{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon_{\theta}}(\boldsymbol{x}_{t},t,\emptyset)}{\sqrt{\bar{\alpha}_{t}}}$ unguided prediction Eq. 10
$M_{t}=\frac{\left\lVert\tilde{\boldsymbol{x}}_{0,t}^{h}-\tilde{\boldsymbol{x}}_{0,t}^{\emptyset}\right\rVert_{2}^{2}}{d}$ divergence Eq. 12
$\boldsymbol{S}_{t}=\tilde{\boldsymbol{x}}_{0,t}^{\emptyset}-\boldsymbol{x}_{0}$ obtain SAM using Eq. 19
end for
$t_{e}=\arg\max_{t}M_{t}$ find the end step
$\boldsymbol{H}=\frac{1}{t_{e}}\sum_{t=1}^{t_{e}}\boldsymbol{S}_{t}$ aggregated SAMs Eq. 20
Obtain the quantile $Q^{*}$ using Algorithm 1 from the main text
[…]
Return predicted pixel-level labels
Algorithm 3 The full segmentation process for a single input $\boldsymbol{x}_{0}$
Appendix D Gradient for Segmentation
We investigate the gradient $\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(h|\boldsymbol{x}_{t}\right)$ as the SAMs for segmentation, while keep other settings unchanged. Here, we use the same BraTS21 testing data, focusing on the unhealthy setup with four configurations: (i) DDIM $\nabla^{2}\log p_{\theta}$: the gradient is directly used as SAMs for segmentation; (ii) DDIM $\nabla^{2}\log p_{\theta}|_{t={t_{e}}}$: only the gradient at the end step is ued for segmentation; (iii) DDIM $C_{t}^{2}\nabla^{2}\log p_{\theta}$: the weighted gradient is used as SAMs; and (iv) DDIM $B_{t}^{2}\left((1+w)\nabla\log p_{\theta}+\Delta\boldsymbol{s}_{t}\right)^{2}$: the original SAMs.
The quantitative results are exhibited in Tab. 3. We note that the performance of other configurations is significantly lower compared to using the original SAMs. The last setup DDIM $C_{t}\nabla\log p_{\theta}$, the weighted gradient, achieved better segmentation results compared to non-weighted gradient SAMs. This is attributed to the weight $C_{t}$, a monotonically increasing function. The SAMs with anomalous regions are more weighted. Also, the error term $\Delta\boldsymbol{s}_{t}$ used in the original SAMs is not considered here, which may alleviate the false detection by the implicit classifier.
Methods DICE IoU AUPRC
DDIM $\nabla^{2}\log p_{\theta}$ 42.7$\pm$0.2 30.6$\pm$0.1 42.0$\pm$0.0
DDIM $\nabla^{2}\log p_{\theta}|_{t={t_{e}}}$ 53.1$\pm$0.0 40.4$\pm$0.0 58.8$\pm$0.0
DDIM $C_{t}^{2}\nabla^{2}\log p_{\theta}$ 57.2$\pm$0.1 45.8$\pm$0.1 70.3$\pm$0.0
DDIM $B_{t}^{2}\left((1+w)\nabla\log p_{\theta}+\Delta\boldsymbol{s}_{t}\right)^{2}$ 61.5$\pm$0.0 51.0$\pm$0.1 75.5$\pm$0.1
Table 3: Segmentation performance on unhealthy samples from BraTS21 dataset using the gradient $\nabla_{\boldsymbol{x}_{t}}\log p_{\theta}\left(h|\boldsymbol{x}_{t}\right)$ as the SAMs. The last setup with * is the original SAMs.
Appendix E More Qualitative Results
We provide more qualitative results for the BraTS21 and ATLAS v2.0 datasets in Fig. 6. It further shows the effectiveness of our method in detecting anomalies and segmenting them. The signal strength of the anomalies is enhanced by the aggregation of SAMs, leading to more accurate segmentation results.
Appendix F Postprocessing for Segmentation
After we obtain the anomaly map, we apply a median filter [12] with kernel size 5 to effectively enhance the performance. Then, we apply the connected component filter to remove the small connected components which is regarded as noise. We apply the same postprocessing to all methods for fair comparison.
Appendix G Discussion and Limitations
All methods showed better results on the BraTS21 dataset than on the ATLAS v2.0 dataset. This disparity arises because DMs are more adept at identifying anomalies that exhibit significant frequency differences, such as tumors on FLAIR MRI, compared to the surrounding healthy tissue. In this case, the difference of healthy and unhealthy distribution is easier to be captured by DMs. In contrast, the ATLAS v2.0 dataset, which consists of T1 MRI, presents more challenging scenarios for anomaly detection due to the subtle frequency differences between healthy and unhealthy regions. During the inference stage, the implicit classifier struggles to accurately capture the anomalous regions, contributing to the less consistent signal of anomalous regions in the SAMs. This inconsistency can lead to the mixing of signals from falsely detected healthy regions, resulting in lower detection accuracy.
Our selection method is specifically designed for the weakly-supervised setting, where unhealthy samples are available for training the guided diffusion model. In unsupervised settings, the unguided diffusion model is typically trained only on healthy samples and evaluated on unhealthy samples. In this scenario, the unguided forward process (UFP) of samples through the diffusion model is not possible, which is a crucial aspect of our method. We leave the exploration of unsupervised settings for future work.
Refer to captionInputGTOurs (DDIM)DDIM clf-freeAno DDPM (S)
Refer to captionInputGTOurs (DDIM)DDIM clf-freeAno DDPM (S)